# Probabilistic Pontryagin’s Maximum Principle for Continuous-Time Model-Based Reinforcement Learning

### v2

@@ Abstract
Without exact knowledge of the true system dynamics, optimal control of non-linear continuous-time systems requires careful treatment of epistemic uncertainty. In this work, we propose a probabilistic extension to Pontryagin’s maximum principle by minimizing the mean Hamiltonian with respect to epistemic uncertainty. We show minimization of the mean Hamiltonian is a necessary optimality condition when optimizing the mean cost, and propose a multiple shooting numerical method scalable to large-scale probabilistic dynamical models, including ensemble neural ordinary differential equations. Comparisons against state-of-the-art methods in online and offline model-based reinforcement learning tasks show that our probabilistic Hamiltonian formulation leads to reduced trial costs in offline settings and achieves competitive performance in online scenarios. By bridging optimal control and reinforcement learning, our approach offers a principled and practical framework for controlling uncertain systems with learned dynamics.
@@ I Introduction
Optimal decision-making under uncertainty is a fundamental challenge in learning-based control, particularly when the system dynamics are unknown and have to be learned from limited and noisy data. Achieving data-efficiency in such settings is crucial, as real-world experiments can incur significant costs. Moreover, effectively handling epistemic uncertainty, which arises from insufficient data, is critical, as it directly influences robustness of control, as well as the exploitation/exploitation balance during learning.
@@
A common approach for decision making under uncertainty involves minimizing the mean cost with respect to the posterior distribution of system dynamics. However, optimizing this mean cost can become challenging in practice, especially when the posterior distribution is multi-modal or has significant variance. Consequently, black-box stochastic optimization methods are typically employed to mitigate local minima, but these ignore the inherent structure of the optimal control problem [1].
@@
Pontryagin’s maximum principle (PMP) provides a different perspective. It characterizes optimal solutions indirectly by describing the necessary conditions of optimality, which implicitly defines the optimal control inputs [2]. As one of the cornerstones in optimal control theory, PMP has driven significant advances, particularly in deterministic control settings [3]. By leveraging the inherent optimality conditions of the control problem, PMP enables computationally efficient approaches to optimization and provides insights into the underlying structure of optimal solutions.
Related work. Principles from optimal control theory serve as foundational tools in learning-based control, providing guidelines for optimal action selection under uncertainty. Several optimal control methodologies have been extended into uncertain or learning-driven contexts, notably the iterative linear quadratic regulator (iLQR) [4] and path integral (PI) control [5]. PMP, originally formulated for deterministic systems, has also seen interest in robust control [6] and, more recently, learning-based settings [7]. An open-loop variant of this was further explored in [8]. However, these prior PMP-based methods are formulated for deterministic system dynamics and do not address epistemic uncertainty arising from limited or noisy data.
@@
Our work closely aligns with Bayesian control paradigms and Bayesian reinforcement learning (RL), where the design of optimal policies explicitly incorporates uncertainty [9]. A seminal work is the probabilistic inference for learning and control (PILCO) framework [10], which learns discrete-time Gaussian process models of system dynamics and efficiently computes policy gradients through closed-form Gaussian approximations. Extensions of Bayesian methods to model predictive control (MPC) have demonstrated efficacy by planning under uncertain system dynamics, leveraging PMP conditions for propagating Gaussian uncertainty over a moving horizon [11]. Model-based planning under uncertainty has also significantly advanced model-based reinforcement learning (MBRL), where methods like probabilistic ensembles with trajectory sampling (PETS) [12] efficiently tackle planning under uncertainty in high-dimensional tasks.
Our work also intersects with recent advances in continuous-time system identification and learning-based control [13], where discrete-time measurements are used to reconstruct the continuous-time dynamics that characterize many physical and biological systems. Extensions into RL have investigated continuous-time policies [14] and efficient exploration [15].
@@
Contribution. Our contributions are threefold. Firstly, we propose a probabilistic extension of classical PMP by minimizing a mean Hamiltonian to explicitly incorporate epistemic uncertainty, and we demonstrate that this provides a necessary optimality condition for control under uncertain dynamics. This probabilistic formulation enhances robustness and performance in scenarios with incomplete or noisy model knowledge. Secondly, we introduce a practical numerical implementation for the mean Hamiltonian approach using a forward multiple shooting method combined with moving horizon optimization, specifically tailored to probabilistic ordinary differential equation (ODE) models. Lastly, we experimentally validate our approach, showing improved or competitive performance compared to state-of-the-art methods in both offline and online RL scenarios, highlighting the value of our method for efficient learning and robust decision-making.
@@ II Problem setting
$\boldsymbol{x}(t)=\boldsymbol{x}_{0}+\int_{t_{0}}^{t_{f}}\boldsymbol{f}\left(\boldsymbol{x}(\tau),\boldsymbol{u}(\tau),\tau\right)\operatorname{d}\!{\tau}$ (II.1)
@@
where $\boldsymbol{x}_{0}\in\mathcal{X}\subset\mathds{R}^{d_{x}}$ is the initial state and $\boldsymbol{u}\colon[t_{0},t_{f}]\to\mathcal{U}\subset\mathds{R}^{d_{u}}$ is the control input. The objective is to find the optimal control input $\boldsymbol{u}^{*}$, that is, the input that minimizes the cost function:
$C(\boldsymbol{u})$ $=\int_{t_{0}}^{t_{f}}L\bigl{(}\boldsymbol{x}(\tau),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\left(\boldsymbol{x}(t_{f})\right)$ (II.2)
@@
where $\Phi(\boldsymbol{x})$ is the termination cost associated with ending in state $\boldsymbol{x}(t_{f})$ at time $t_{f}$ and $L(\boldsymbol{x},\boldsymbol{u})$ is the state cost associated with taking action $\boldsymbol{u}$ in state $\boldsymbol{x}$. Both the termination cost and state cost are assumed differentiable with respect to their inputs. We are interested in solving the following optimal control problem:
$\boldsymbol{u}^{*}(t)$ :-$\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}C(\boldsymbol{u}):$ (II.3)
[…]
$\boldsymbol{x}(t_{0})$ $=\boldsymbol{x}_{0}\enspace.$
@@
Although the true function $\boldsymbol{f}$ is unknown, it can be evaluated by interacting with the system and collecting data over the course of $N$ trials with time horizon $t_{f}$. Let $\mathcal{D}_{n}\coloneq\{(t^{n}_{i},\boldsymbol{y}^{n}_{i},\boldsymbol{u}^{n}_{i})\}$ denote the data collected in trial $n$. Here, $t^{n}_{i}\in[t_{0},t_{f}]$ is the time point corresponding to observation $i$ in trial $n$, $\boldsymbol{x}_{i}^{n}\coloneq\boldsymbol{x}(t^{n}_{i})$ is the true system state at time $t^{n}_{i}$ in trial $n$, $\boldsymbol{y}^{n}_{i}\sim\mathcal{N}(\boldsymbol{x}_{i}^{n},\sigma_{n}^{2}\boldsymbol{I})$ is a noisy measurement of the state with i.i.d. Gaussian noise, and $\boldsymbol{u}^{n}_{i}$ is the applied control input corresponding to observation $i$ for trial $n$. It is assumed that the true control function can be reconstructed for $t\in[t_{0},t_{f}]$ using interpolation techniques.
@@
In MBRL, a dynamical model is learned based on previous interactions with the system. In the online variant, this dataset consists initially of only a single trial generated by a random controller. Then iteratively, a model is trained on the available data, the optimal action is determined using the model, and the newly observed trajectory is added to the data.
II-A Probabilistic dynamical models
@@
When learning a dynamical model from limited data, it is crucial to employ efficient exploration strategies that separate epistemic uncertainty, associated with reducible uncertainty due to lack of knowledge of the true system, from aleatoric uncertainty, associated with irreducible uncertainty due to inherent randomness such as sensor or process noise. A general approach for quantifying uncertainty is to assign a prior distribution over the parameters and a likelihood to the possibly noisy observations, and use Bayes’ theorem to marginalize over the uncertainty in model parameters:
$\boldsymbol{\theta}$ $\sim p(\boldsymbol{\theta})$ (II.4)
$p(\mathcal{D}\mid\boldsymbol{\theta})$ $=\prod_{i=1}^{T}\mathcal{N}\left(\boldsymbol{y}_{i}\mid\boldsymbol{x}(t_{i};\boldsymbol{\theta}),\sigma^{2}\boldsymbol{I}\right)$ (II.5)
$p(\boldsymbol{\theta}\mid\mathcal{D})$ $\propto p(\mathcal{D}\mid\boldsymbol{\theta})\>p(\boldsymbol{\theta})$ (II.6)
@@
where $T$ is the number of discrete-time observations throughout $[t_{0},t_{f}]$ and $\boldsymbol{\theta}\in\Theta=\mathbb{R}^{p}$ are the parameters. For a given realization $\boldsymbol{\theta}$, the following continuous-time model is defined:
$\dot{\boldsymbol{x}}(t;\boldsymbol{\theta})$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{u}(t),t;\boldsymbol{\theta})$ (II.7)
@@
where $\boldsymbol{f}$ is differentiable in its first two arguments. The posterior distribution can then be estimated using Bayesian sampling methods, such as Hamiltonian Monte Carlo (HMC).
@@ II-B Mean cost objective
A common objective in MBRL is the expected cost over the parameter distribution of the dynamical model, which reflects the epistemic uncertainty:
$\bar{C}(\boldsymbol{u})\coloneq\mathbb{E}_{\boldsymbol{\theta}}\left[C(\boldsymbol{u};\boldsymbol{\theta})\right]$ (II.8)
@@
$C(\boldsymbol{u};\boldsymbol{\theta})$ $=\int_{0}^{t_{f}}L\bigl{(}\boldsymbol{x}(\tau;\boldsymbol{\theta}),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta})\bigr{)}$
and $\mathbb{E}_{\boldsymbol{\theta}}$ is shorthand for $\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})}$. The cost functional $C(\boldsymbol{u};\boldsymbol{\theta})$ denotes the cost for realization $\boldsymbol{\theta}$. This results in the optimal control problem:
$\boldsymbol{u}^{*}(t)$ :-$\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}\bar{C}(\boldsymbol{u}):$ (II.9)
[…]
$\boldsymbol{x}(t_{0};\boldsymbol{\theta})$ $=\boldsymbol{x}_{0}\enspace.$
@@ III-A Pontryagin’s maximum principle
The mean cost objective can be optimized via the direct method or via the indirect method. In the former, the control function is discretized and optimized as a real-valued vector, whereas in the latter, first the necessary optimality conditions for the continuous-time optimal control problem are described using PMP and then the control function is discretized. A key function is the Hamiltonian:
$\mathcal{H}\bigl{(}\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t\bigr{)}\coloneq L\bigl{(}\boldsymbol{x},\boldsymbol{u}\bigr{)}+\boldsymbol{\lambda}^{\top}\boldsymbol{f}\bigl{(}\boldsymbol{x},\boldsymbol{u},t\bigr{)}$
@@
$\dot{\boldsymbol{x}}(t)$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t),\boldsymbol{u}(t),t\bigr{)}$ (III.1)
$\dot{\boldsymbol{\lambda}}(t)$ $=-\nabla_{\boldsymbol{x}}\mathcal{H}\bigl{(}\boldsymbol{x}(t),\boldsymbol{\lambda}(t),\boldsymbol{u}(t),t\bigr{)}$ (III.2)
$\boldsymbol{\lambda}(t_{f})$ $=\nabla\Phi\bigl{(}\boldsymbol{x}(t_{f})\bigr{)}$ (III.3)
for $t\in[t_{0},t_{f}]$, where the notation $(\nabla_{y}f)_{i,j}\coloneq\partial f_{j}/{\partial y_{i}}$ denotes Jacobians, in this case $\nabla_{\boldsymbol{x}}\mathcal{H}=\left(\frac{\partial\mathcal{H}}{\partial x_{1}},\dots,\frac{\partial\mathcal{H}}{\partial x_{d_{x}}}\right)^{\top}$ and $\nabla\Phi=\left(\frac{\partial\Phi}{\partial x_{1}},\dots,\frac{\partial\Phi}{\partial x_{d_{x}}}\right)^{\top}$. Equation (III.3) is commonly known as the transversality condition, and provides a boundary condition for the co-states. If $\boldsymbol{u}^{*}(t)$ is an optimal solution to $C(\boldsymbol{u})$, the following necessary condition of optimality is derived via PMP:
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}\left(\boldsymbol{x}^{*}(t),\boldsymbol{\lambda}^{*}(t),\boldsymbol{u},t\right)$ (III.4)
where $\boldsymbol{x}^{*}(t)$ and $\boldsymbol{\lambda}^{*}(t)$ satisfy conditions (III.1) – (III.3). We will refer to equation (III.4) as the minimum principle. The indirect approach provides candidate solutions for the optimal control $\boldsymbol{u}^{*}(t)$ that satisfy the necessary optimality conditions for the cost functional. Next, it is detailed how these conditions extend to the probabilistic setup.

### v3

# Optimal Control of Probabilistic Dynamics Models via Mean Hamiltonian Minimization

@@ Abstract
Without exact knowledge of the true system dynamics, optimal control of non-linear continuous-time systems requires careful treatment under epistemic uncertainty. In this work, we translate a probabilistic interpretation of the Pontryagin maximum principle to the challenge of optimal control with learned probabilistic dynamics models. Our framework provides a principled treatment of epistemic uncertainty by minimizing the mean Hamiltonian with respect to a posterior distribution over the system dynamics.
We propose a multiple shooting numerical method that leverages mean Hamiltonian minimization and is scalable to large-scale probabilistic dynamics models, including ensemble neural ordinary differential equations. Comparisons against other baselines in online and offline model-based reinforcement learning tasks show that our probabilistic Hamiltonian approach leads to reduced trial costs in offline settings and achieves competitive performance in online scenarios. By bridging optimal control and reinforcement learning, our approach offers a principled and practical framework for controlling uncertain systems with learned dynamics.
@@ I Introduction
Optimal decision-making under uncertainty is a fundamental challenge in learning-based control, especially when the system dynamics are unknown and must be inferred from limited, noisy data. In such settings it is crucial to achieve high data-efficiency, as real-world experiments often come with high costs or risks. To address this, one has to effectively handle epistemic uncertainty, arising from incomplete data, which critically shapes both the robustness of the controller and the balance between exploration and exploitation in learning.
@@
A common approach for decision making under uncertainty involves minimizing the mean cost over the posterior distribution of system dynamics. However, this optimization can become challenging when the posterior distribution is multi-modal or exhibits significant variance. While black-box stochastic optimization methods are often employed to avoid local minima, they overlook the inherent structure of the optimal control problem [1].
@@
Pontryagin’s maximum principle (PMP) offers a compelling alternative. By defining necessary conditions for optimality, PMP indirectly characterizes optimal control inputs, providing a powerful lens into the structure of optimal solutions [2]. As a cornerstone of optimal control theory, PMP has driven significant advances in deterministic and stochastic control settings [3, 4]. Its ability to leverage the inherent structure of the optimal control problem enables computationally efficient optimization and insights into optimal behavior.
Figure 1: Optimal control of an uncertain dynamical system using the mean Hamiltonian minimization principle. (Left) The Hamiltonian vector field for a system with dynamics depicting state and co-states of the dynamics function $f_{\theta}(x,u)=\theta\sin(x)+u$. The thin lines show posterior trajectories sampled from the dynamics belief $p(\theta)$, while the bold black line shows the nominal solution of the control function that minimizes the mean Hamiltonian. (Middle & Right) The evolution of the state $x(t)$, co-state $\lambda(t)$, optimal control $u^{*}(t)$, and the mean Hamiltonian $\mathbb{E}_{\theta}[\mathcal{H}_{\theta}]$. The optimal control function, derived from minimizing the mean Hamiltonian, successfully steers the system to the origin. As a necessary condition for optimality, the mean Hamiltonian is constant under the optimal control.
Related work. Optimal control theory provides foundational principles for learning-based control, guiding optimal action selection under uncertainty. Techniques like the iterative Linear Quadratic Regulator (iLQR) [5] and path integral (PI) control [6] have been widely adopted to design optimal policies in uncertain or learning-driven settings. Originally developed for deterministic systems, PMP has also been extended to stochastic [4], robust [7], and, more recently, learning-based control [8, 9]. While these PMP-based methods primarily address deterministic systems or those with aleatoric uncertainty, recent theoretical advances have established optimality conditions for systems with epistemic uncertainty in their dynamics. In [10], conditions are derived for minimizing average terminal costs in general systems with parametric uncertainty, while in [11] similar conditions are developed for control-affine ensemble systems. Our work builds on this foundation by operationalizing the principle of mean Hamiltonian minimization for learned probabilistic dynamics.
@@
Our approach closely aligns with Bayesian control and reinforcement learning (RL) paradigms, where optimal policies explicitly account for uncertainty [12, 13]. The probabilistic inference for learning and control (PILCO) framework [14] exemplifies this by learning a Gaussian process (GP) dynamics model and computing policy gradients via closed-form Gaussian approximations. Similarly, Bayesian extensions to model predictive control (MPC) have shown efficacy by planning under uncertain dynamics, using PMP conditions to propagate Gaussian uncertainty over a moving horizon [15].
In deep model-based reinforcement learning (MBRL), methods like probabilistic ensembles with trajectory sampling (PETS) [16] have advanced high-dimensional task solving through uncertainty-aware planning. Our work bridges these advances with continuous-time RL, where discrete-time measurements inform continuous-time dynamics and control [17, 18]. By integrating PMP with these Bayesian and continuous-time approaches, we propose a novel framework for optimal planning under epistemic uncertainty, tailored to ensemble-based RL, offering a robust and principled approach to controlling complex, uncertain systems.
@@
Contribution. Our contributions are threefold. First, we formulate the optimal control problem for learned probabilistic dynamics models within the PMP framework, introducing a principled approach to planning based on mean Hamiltonian minimization that explicitly addresses epistemic uncertainty. Second, we develop a practical numerical method combining forward multiple shooting with moving horizon optimization, tailored to probabilistic ordinary differential equation (ODE) models. Third, we validate our approach experimentally, demonstrating improved or competitive performance compared to state-of-the-art methods in both offline and online RL scenarios, highlighting its potential for efficient learning and robust decision-making.
The application of our framework to a representative optimal control problem with uncertain dynamics is illustrated in Figure 1, which visualizes the full probabilistic solution, from the state-costate trajectories to the final optimal control.
@@ II Problem setting
$\boldsymbol{x}(t)=\boldsymbol{x}_{0}+\int_{t_{0}}^{t_{f}}\boldsymbol{f}\left(\boldsymbol{x}(\tau),\boldsymbol{u}(\tau),\tau\right)\operatorname{d}\!{\tau},$ (II.1)
@@
where $\boldsymbol{x}_{0}\in\mathcal{X}\subset\mathds{R}^{d_{x}}$ is the initial state, $\boldsymbol{u}\colon[t_{0},t_{f}]\to\mathcal{U}\subset\mathds{R}^{d_{u}}$ is the control input, and $d_{x}$ and $d_{u}$ denote the dimensions of the state space and the control input. The objective is to find the optimal control input $\boldsymbol{u}^{*}$, that is, the input that minimizes the cost function:
$C(\boldsymbol{u})$ $\coloneq\int_{t_{0}}^{t_{f}}L\bigl{(}\boldsymbol{x}(\tau),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\left(\boldsymbol{x}(t_{f})\right),$ (II.2)
@@
where $\Phi(\boldsymbol{x})$ is the terminal cost associated with ending in state $\boldsymbol{x}(t_{f})$ at time $t_{f}$ and $L(\boldsymbol{x},\boldsymbol{u})$ is the state cost associated with taking action $\boldsymbol{u}$ in state $\boldsymbol{x}$. Both the terminal cost and state cost are assumed differentiable with respect to their inputs. We are interested in solving the following optimal control problem:
$\boldsymbol{u}^{*}(t)$ $\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}C(\boldsymbol{u})$ (II.3)
$\text{s.t.}\quad\dot{\boldsymbol{x}}(t)$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t),\boldsymbol{u}(t),t\bigr{)},\quad\text{for }t\in[t_{0},t_{f}]\enspace$
$\boldsymbol{x}(t_{0})$ $=\boldsymbol{x}_{0}.$
@@
Although the true function $\boldsymbol{f}$ is unknown, it can be evaluated by interacting with the system and collecting data over the course of $N$ trials over the time horizon $t_{f}$. Let $\mathcal{D}_{n}\coloneq\{(t^{n}_{i},\boldsymbol{y}^{n}_{i},\boldsymbol{u}^{n}_{i})\}$ denote the data collected in trial $n\in[1,N]$. Here, $t^{n}_{i}\in[t_{0},t_{f}]$ is the time point corresponding to observation $i$ in trial $n$, $\boldsymbol{x}_{i}^{n}\coloneq\boldsymbol{x}(t^{n}_{i})$ is the true system state at time $t^{n}_{i}$ in trial $n$, $\boldsymbol{y}^{n}_{i}\sim\mathcal{N}(\boldsymbol{x}_{i}^{n},\sigma_{n}^{2}\boldsymbol{I})$ is a noisy measurement of the state with i.i.d. Gaussian noise with variance $\sigma_{n}^{2}$, and $\boldsymbol{u}^{n}_{i}$ is the applied control input corresponding to observation $i$ for trial $n$, assumed to be exactly known. It is assumed that the true control function can be reconstructed for $t\in[t_{0},t_{f}]$ using interpolation techniques.
@@
In MBRL, a dynamics model is learned based on previous interactions with the system. In the online variant, this dataset consists initially of only a single trial generated by a random controller. Then iteratively, a model is trained on the available data, the optimal action is determined using the model, and the newly observed trajectory is added to the data.
II-A Probabilistic dynamics models
@@
When learning a dynamics model from limited data, it is crucial to employ efficient exploration strategies that separate epistemic uncertainty, associated with reducible uncertainty due to lack of knowledge of the true system, from aleatoric uncertainty, associated with irreducible uncertainty due to inherent randomness such as sensor or process noise. A general approach for quantifying uncertainty is to assign a prior distribution over the parameters and a likelihood to the possibly noisy observations, and use Bayes’ theorem to obtain the posterior distribution over model parameters:
$\boldsymbol{\theta}$ $\sim p(\boldsymbol{\theta}),$ (II.4)
$p(\mathcal{D}\mid\boldsymbol{\theta})$ $=\prod_{i=1}^{T}\mathcal{N}\left(\boldsymbol{y}_{i}\mid\boldsymbol{x}_{\boldsymbol{\theta}}(t_{i}),\sigma^{2}\boldsymbol{I}\right),$ (II.5)
$p(\boldsymbol{\theta}\mid\mathcal{D})$ $\propto p(\mathcal{D}\mid\boldsymbol{\theta})\>p(\boldsymbol{\theta}).$ (II.6)
@@
Here, $T$ is the number of discrete-time observations throughout $t_{0},t_{f}]$ and $\boldsymbol{\theta}\in\Theta=\mathbb{R}^{p}$ are the parameters. For a given realization $\boldsymbol{\theta}$, the following continuous-time model is defined:
$\dot{\boldsymbol{x}}_{\boldsymbol{\theta}}(t)$ $=\boldsymbol{f}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{u}(t),t),$ (II.7)
@@
where $\boldsymbol{f}_{\boldsymbol{\theta}}$ is differentiable in its first two arguments. The posterior distribution can then be estimated using Bayesian sampling methods such as Hamiltonian Monte Carlo (HMC) [19].
@@ II-B Mean cost objective
A common objective in MBRL is the expected cost over the parameter distribution of the dynamics model which reflects the epistemic uncertainty:
$\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})\coloneq\mathbb{E}_{\boldsymbol{\theta}}\left[C_{\boldsymbol{\theta}}(\boldsymbol{u})\right],$ (II.8)
@@
$C_{\boldsymbol{\theta}}(\boldsymbol{u})$ $\coloneq\int_{0}^{t_{f}}L\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(\tau),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t_{f})\bigr{)},$
and $\mathbb{E}_{\boldsymbol{\theta}}$ is shorthand for $\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})}$. This results in the optimal control problem:
$\boldsymbol{u}^{*}(t)$ $\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})$ (II.9)
$\text{s.t.}\quad\dot{\boldsymbol{x}}_{\boldsymbol{\theta}}(t)$ $=\boldsymbol{f}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{u}(t),t\bigr{)},\quad\text{for }t\in[t_{0},t_{f}]\enspace$
$\boldsymbol{x}_{\boldsymbol{\theta}}(t_{0})$ $=\boldsymbol{x}_{0}.$
@@ III-A Pontryagin’s maximum principle
The mean cost objective $\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})$ can be optimized via the direct method or via the indirect method. In the former, the control function is discretized and optimized as a real-valued vector, whereas in the latter, first the necessary optimality conditions for the continuous-time optimal control problem are described using PMP and then the control function is discretized. A key function is the Hamiltonian:
$\mathcal{H}\bigl{(}\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t\bigr{)}\coloneq L\bigl{(}\boldsymbol{x},\boldsymbol{u}\bigr{)}+\boldsymbol{\lambda}^{\top}\boldsymbol{f}\bigl{(}\boldsymbol{x},\boldsymbol{u},t\bigr{)},$
@@
$\dot{\boldsymbol{x}}(t)$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t),\boldsymbol{u}(t),t\bigr{)},$ (III.1)
$\dot{\boldsymbol{\lambda}}(t)$ $=-\nabla_{\boldsymbol{x}}\mathcal{H}\bigl{(}\boldsymbol{x}(t),\boldsymbol{\lambda}(t),\boldsymbol{u}(t),t\bigr{)},$ (III.2)
$\boldsymbol{\lambda}(t_{f})$ $=\nabla\Phi\bigl{(}\boldsymbol{x}(t_{f})\bigr{)},$ (III.3)
for $t\in[t_{0},t_{f}]$, where the notation $(\nabla_{y}f)_{i,j}\coloneq\partial f_{j}/{\partial y_{i}}$ denotes Jacobians, in this case $\nabla_{\boldsymbol{x}}\mathcal{H}=(\frac{\partial\mathcal{H}}{\partial x_{1}},\dots,\frac{\partial\mathcal{H}}{\partial x_{d_{x}}})^{\top}$ and $\nabla\Phi=(\frac{\partial\Phi}{\partial x_{1}},\dots,\frac{\partial\Phi}{\partial x_{d_{x}}})^{\top}$. Equation (III.3) is commonly known as the transversality condition, and provides a boundary condition for the co-states. If $\boldsymbol{u}^{*}(t)$ is an optimal solution to $C(\boldsymbol{u})$, the following necessary condition of optimality is derived via PMP:
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}\left(\boldsymbol{x}^{*}(t),\boldsymbol{\lambda}^{*}(t),\boldsymbol{u},t\right),$ (III.4)
where $\boldsymbol{x}^{*}(t)$ and $\boldsymbol{\lambda}^{*}(t)$ satisfy (III.1) – (III.3). We refer to (III.4) as the minimum principle. The indirect approach thus provides candidate solutions for the optimal control $\boldsymbol{u}^{*}(t)$ that satisfy the necessary optimality conditions for the cost functional. Next, we extend these conditions to the probabilistic setup.